Long Questions (90%) (1) Simple Issues with Measuring GDP Suppose the economy consists of 2 people. Person 1 produces 1 unit of goods each period, which is split between the 2 pcople. (a) What is per-period real aggregate GDP? (b) What is per-capita real GDP in each period? (c) Suppose now that 20 percent of each periods production is saved for later (i.e. not consumed). What is aggregate GDP? What is per-capita GDP? Assume the saved goods ncither go bad nor carn interest (i.c. the rate of return on saving equals zero
Answer #1

Solution(s):

Real GDP is 1 unit of goods which is produced by product method. By expenditure method=0.5+0.5=1 unit GDP includes goods produced

B 1/2=0.5.per capita means per individual. There are two people and output is 1.Real gdp=output/population

C Aggregate GDP remains same. Consumption is equal to invest in national income accounting. Alternatively saving comes from income. And by income method gdp remains 1.